OpenStudy (anonymous):

what is the volume of water at 1 ATM evaporated from 0.1 ml of water?

OpenStudy (anonymous):

You mean volume of water vapour ? Considering that water vapours will behave like ideal gas, we can apply ideal gas law to them. PV = nRT P = 1 atm =1.013 * 10^5 N/m2 R = 8.314 J/K/mol To find n, number of moles , we notice n=mass of gas(m) / molar mass of gas(M) the volume of water is 0.1ml density of water is 1g/ml so m = 0.1g M = 18 g Hence we can get the no of moles Your question does not provide value of temperature at which it evaporates. It may be room temperature (298K) V is the volume of water vapour which we can get from the above eqn.